# Spatiotemporal-Augmented Graph Neural Networks for Human Mobility Simulation

### v1

@@ 4.1.1 Location Graphs
Let $\mathcal{P}=\{p_{0},p_{1},\cdots,p_{N}\}$ be the set of locations. The spatiotemporal correspondences among locations $p_{i},p_{j}\in\mathcal{P}$ can be well modeled by their correspondence scores $\epsilon(p_{i},p_{j})$ with a correspondence function $\epsilon(\cdot,\cdot)$, resulting in a location-location adjacency graph $G=(\mathcal{P},\mathcal{E})$. The edge set $\mathcal{E}$ of the graph contains pairwise correspondences of locations. For example, if the function $\epsilon(\cdot,\cdot)$ measures the geographical proximity of locations, the constructed graph $G$ represents a $K$-nearest neighbor spatial graph of all locations. Based on the spatial graph, our model will generate human mobility trajectories in consideration of geographical effects.
Therefore, to go beyond the geographical effects, we firstly build a Spatial Distance Graph (SDG) and Temporal Transition Graph (TTG) based on the observed human mobility trajectories, which are constructed from the spatial and temporal perspectives, respectively. Secondly, we propose a SpatioTemporal Graph (STG) by measuring the Wasserstein distance [60] of location visit distributions to capture the spatiotemporal effects, which combines the spatial and temporal dynamics of locations simultaneously.
Spatial Distance Graph. The proposed SDG is constructed based on the spatial proximity of a location pair $(p_{i},p_{j})$, which indicates the spatial cooperation effects in human trajectories. Let $w_{ij}$ be the spatial proximity score of $\epsilon(p_{i},p_{j})$, we have to build the Spatial Distance Graph (SDG) out of the Cartesian score set of $\mathcal{P}\times\mathcal{P}$, written as
@@
which remains the top-$k$ neighbors from all locations. The spatial proximity function is usually implemented by Euclidean distance for simplicity.
Temporal Transition Graph. The SDG can reveal the spatial proximity of locations, but ignores the mobility patterns of human trajectories. We further propose the Temporal Transition Graph (TTG) to encode the mobility patterns based on the partial observed human trajectories. Let $s=[\cdots,\tau_{l},\tau_{l+1},\cdots]$ be a human trajectory and $\tau_{l}=(p_{i},t_{l}),\tau_{l+1}=(p_{j},t_{l+1})$. We record the proximity score between $p_{i}$ and $p_{j}$ as $\epsilon(p_{i},p_{j})=1$ and obtain the summed scores by $w_{ij}=\sum_{(pi,p_{j})\in\mathcal{S}}\epsilon(p_{i},p_{j})$. The graph is thus formulated as
@@
The obtained TTG can detailedly describe the mobility patterns from the temporal transition perspective.
SptioTemporal Graph. The SDG and TTG describe the location relationships from the spatial and temporal perspectives, respectively, thus ignore the other aspect to some extent. We further characterize the location functionality with a SpatioTemporal Graph (STG) by its visit distribution over time, which can describe the spatial proximity with the spatial cooperation effects and the temporal proximity with the visit distribution similarity. For instance, cafes often open from morning to evening, while bars keep open till midnight. The visit distribution $F_{p_{i}}$ of a location $p_{i}$ is calculating its normalized visit counts in the observed records $\mathcal{S}_{p_{i}}$ by discretizing the timestamps into $T$ time slots: $F_{p_{i}}(t)=\frac{1}{|\mathcal{S}_{p_{i}}|}\sum_{t_{k}\in\mathcal{S}_{p_{i}}}\mathbbm{1}{[t_{k}=t]}$. The obtained visit distribution $F_{i}$ can well describe the location from both the spatial and the temporal aspects. On the one hand, neighborhood locations often share similar visit distributions based on geographical effects, which have been widely adopted in Location-based Services. On the other hand, locations far apart can also share similar visit distributions since they provide peer functionalities like eating or drinking. Thus, the Wasserstein distance [60] is employed here to measure the distances of the visit distributions $F_{\mathcal{P}}$ of locations to differentiate their individual functionalities, defined as follows:
@@
where $\pi(x,y)$ is a joint distribution of $F_{p_{i}}$ and $F_{p_{j}}$. The proximity score can be simply defined as $\epsilon(p_{i},p_{j})=1-d(p_{i},p_{j})\in[0,1]$. Let $w_{ij}=\epsilon(p_{i},p_{J})$, the graph can be formulated as
@@
The proposed STG can capture the spatiotemporal dynamics of locations in a distribution way.
Overall, the above multi-channel location graphs are denoted by $\mathcal{G}=\{G_{\mathrm{SDG}},G_{\mathrm{TTG}},G_{\mathrm{STG}}\}$. The connection edges of a location pair $(p_{i},p_{j})\in G$ are naturally weighted with their proximity scores, indicating the spatiotemporal correspondences given by different measurements. However, these scores are not aligned with the human mobility simulation task, which can cause the inconsistent problems between the spatiotemporal graphs and the simulation task. To bridge the gap between the graphs and the task, we alternatively propose a Vanilla version based on the Weighted version, written as
@@
The Vanilla graph regards all neighbors as the same, benefiting for the model optimization without the computed weights of the multi-channel graphs.
@@ 4.1.2 Location Embeddings
Firstly, we implement the features of the location set $\mathcal{P}$ at the input layer by a learnable embedding matrix with end-to-end optimization, following the common training paradigm of human mobility simulation. Secondly, to adaptively aggregate the useful information in terms of the graph heterogeneity, we adopt the attention layer of graphs [45] to perform multi-channel attention over all kinds of spatiotemporal graphs, namely $\mathcal{G}=\{G_{\mathrm{SDG}},G_{\mathrm{TTG}},G_{\mathrm{STG}}\}$. Given location embeddings $\mathbf{H}^{l-1}_{G}$ from the previous $l-1$-th layer (initialized embeddings at the input layer), our spatiotemporal attention layer computes the node embeddings in a multi-head manner by
@@
For simplicity, the $l$-th layer location embeddings are summed over all graphs: $\mathbf{H}^{l}=\sum_{G\in\mathcal{G}}\mathbf{H}^{l}_{G}$. The final location embeddings are denoted by $\mathbf{H}_{\mathcal{P}}$.
@@ 4.2 Decision Generator Module
The multi-channel embedding module provides location embeddings via our dedicated spatiotemporal graphs, enhancing the representation quality used in the following decision generator module. The decision generator module predicts the next location based on the observed partial mobility trajectory, which is also generated by the decision generator sequentially. Besides the common exploration branch of GANs, we further propose the dwell branch to cope with the varying durations of locations as shown in Figure 1 (b), which also reduces the optimization difficulty to some extent. Since the policy discriminator will not give classification results until the trajectory sequence is finished, we adopt the Monte-Carlo strategy to iteratively predict the next location until the required length.
@@ 4.2.1 Exploration Branch
Suppose that a partial mobility trajectory $s_{1:l}=\{(p_{1},t_{1}),(p_{2},t_{2}),\cdots,(p_{l},t_{l})\}$ is observed, the exploration branch aims to predict the next location $\hat{p}_{l+1}$ that maximizes the decision reward given by the policy discriminator module, where $t_{l+1}$ is assumed uniformly spaced with $t_{l}$ for simplicity. Given a bunch of location embeddings $\mathbf{H}_{\mathcal{P}}$ from the multi-channel embedding module, we firstly fetch the corresponding input embeddings of the trajectory, denoted by $\mathbf{h}_{s}=\{\mathbf{h}_{p_{0}},\mathbf{h}_{p_{1}},\cdots,\mathbf{h}_{p_{l}}\}$. To facilitate the sequence modeling and training efficiency, we adopt the Gated Recurrent Unit [61] as our sequence model, written as
@@
where the hidden state $z_{l-1}$ is initialized as zero vectors in the first step. The obtained sequence representation $\mathbf{h}_{l}$ is fed into a linear layer to predict the probability of the next locations by
@@
Then the outputs of the exploration branch $\hat{p}_{l+1}$ together with the outputs of the dwell branch determine the next location iteratively until the sequence is finished.
@@ 4.2.2 Dwell Branch
Motivated by the varying durations of locations in mobility trajectories, we specially design a dwell branch to predict the probability of staying at the previous location, which can inform the model to dwell or explore. Intuitively, the duration of a specific location is determined by the spatiotemporal context, and the cumulative duration of the trajectory, where the former can be well captured by the exploration branch and the latter lacks specific modeling. For a highly repetitive trajectory dataset, the optimization goal will be overwhelmed by the repetitive locations, leading to the ignorance of diverse behaviors. Therefore, it is necessary to build the dwell branch to predict whether to stay at the previous location to alleviate the learning difficulties of the exploration branch. Taking the sequence representation $\mathbf{h}_{l}$ for the dwell branch classification is straightforward. However, it ignores the saturability of human mobility behaviors [1] that the durations of locations are usually upper-bounded to some extent. We thus combine the sigmoid function with an exponentially decaying coefficient to reduce the effects of a long-duration location, written as
@@
Finally, the decision generator module balances the location predictions $\hat{p}_{l+1}$ of the exploration branch and the dwell predictions $\hat{y}_{l+1}^{d}$ by sampling from the output distributions, written as
@@ 4.3 Policy Discriminator Module
By generating and sampling iteratively in the decision generator module, a set of generated trajectories $\mathcal{T}_{G}$ are fed into the policy discriminator module $D_{\phi}$. Since a powerful discriminator might hinder the optimization of the generator $G_{\theta}$, we build a simple yet efficient discriminator $D_{\phi}$ that contains an embedding matrix of locations and classifies the trajectory sequence as real data or fake data based on a GRU layer, written as
@@
where $\mathcal{T}_{R}$ and $\mathcal{T}_{G}$ are real and generated trajectories, respectively. The step-by-step predictions of the decision generator module $G_{\theta}$ receives the policy gradients from the discriminator $D_{\phi}$ by unfolding the classification reward recursively following the REINFORCE algorithm [62], written as
@@ 5 Experiment
TABLE II: Performance of the proposed STAR framework and baselines in terms of JSD for human mobility trajectory simulation. A lower JSD value indicates a better performance. Bold and underline means the best and the second-best results.
@@
Markov[10] 0.1256 $\underline{0.3742}$ 0.0037 0.2667 0.1119 0.0668 0.1379 0.4246 0.0077 0.2615 0.0413 0.0659
IOHMM[11] 0.4179 0.6022 0.0046 0.1071 0.6931 0.0750 0.3368 0.5438 0.0059 0.1732 0.6931 0.0577
LSTM[63] $\underline{0.1239}$ 0.3862 0.0014 $\underline{0.0776}$ $\underline{0.0772}$ 0.0728 0.1334 0.4479 0.0019 $\underline{0.0415}$ 0.0350 0.0615
DeepMove[12] 0.4068 0.5851 0.0064 0.3139 0.4030 0.0593 0.3349 0.5374 $\underline{0.0010}$ 0.0478 0.2926 0.0534
GAN[15] 0.4405 0.6298 0.0026 0.0839 0.2418 0.0364 0.3644 0.5653 0.0046 0.1881 0.1599 0.0557
SeqGAN[33] 0.1774 0.4374 0.0043 0.0803 0.1159 0.0528 $\underline{0.1260}$ $\underline{0.4193}$ 0.0011 0.0452 $\underline{0.0267}$ 0.0591
MoveSim[13] 0.4571 0.5977 0.0013 0.0885 0.2474 0.0624 0.3459 0.5511 0.0036 0.1655 0.0777 $\underline{0.0476}$
CGE[14] 0.4154 0.6338 0.1101 0.3925 0.3172 $\underline{0.0476}$ 0.4151 0.5816 0.0041 0.1792 0.1820 0.0457
STAR (Ours) 0.1105 0.3636 $\underline{0.0011}$ 0.0580 0.0359 0.0624 0.1196 0.4178 0.0005 0.0402 0.0259 0.0693
@@
Markov[10] $\underline{0.0083}$ 0.0555 0.0136 0.1729 0.1229 0.0694 0.0078 0.0409 0.0060 0.1427 0.2505 0.0797
IOHMM[11] 0.1050 0.0610 0.0137 0.2832 0.1639 0.0733 0.0447 0.0456 0.0053 0.1729 0.1839 0.0693
LSTM[63] 0.0127 0.0208 0.0192 0.4412 0.3621 0.0626 $\underline{0.0066}$ $\underline{0.0097}$ 0.0777 0.6457 0.2965 0.0596
DeepMove[12] 0.1098 0.0528 0.0046 0.1592 0.3814 0.0532 0.0365 0.0279 0.0018 $\underline{0.0405}$ 0.2931 $\underline{0.0459}$
GAN[15] 0.1073 0.0735 0.0098 0.2511 0.2158 0.0755 0.0494 0.0438 0.0042 0.1608 0.3325 0.0577
SeqGAN[33] 0.0123 0.0207 $\underline{0.0036}$ 0.0434 $\underline{0.0307}$ 0.0659 0.0106 0.0106 0.0028 0.0696 $\underline{0.0294}$ 0.0451
MoveSim[13] 0.0235 $\underline{0.0179}$ 0.0457 0.4915 0.1252 $\underline{0.0523}$ 0.0096 0.0099 $\underline{0.0016}$ 0.5643 0.2459 0.0487
CGE[14] 0.0328 0.6931 0.0062 0.2322 0.2170 0.0659 0.1798 0.1008 0.0785 0.5164 0.3254 0.0581
STAR (Ours) 0.0049 0.0152 0.0023 $\underline{0.0497}$ 0.0294 0.0485 0.0062 0.0095 0.0014 0.0371 0.0287 0.0485

### v3

@@ 4.1.1 Location Graphs
Let $\mathcal{P}=\{p_{0},p_{1},\cdots,p_{N}\}$ be the set of locations. The spatiotemporal correspondences among locations $p_{i},p_{j}\in\mathcal{P}$ can be well modeled by their correspondence scores $\epsilon(p_{i},p_{j})$ with a correspondence function $\epsilon(\cdot,\cdot)$, resulting in a location-location adjacency graph $G=(\mathcal{P},\mathcal{E})$. The edge set $\mathcal{E}$ of the graph contains pairwise correspondences between locations. For example, if the function $\epsilon(\cdot,\cdot)$ measures the geographical proximity of locations, the constructed graph $G$ represents a $K$-nearest neighbor spatial graph of all locations. Based on the spatial graph, our model will generate human mobility trajectories in consideration of geographical effects. Next, we describe the construction of Spatial Distance Graph, Temporal Transition Graph and SpatioTemporal Graph respectively.
Spatial Distance Graph. The proposed SDG is constructed based on the spatial proximity of a location pair $(p_{i},p_{j})$, which indicates the spatial cooperation effects in human trajectories. Let $w_{ij}$ be the spatial proximity score of $\epsilon(p_{i},p_{j})$, we have to build the Spatial Distance Graph out of the Cartesian score set of $\mathcal{P}\times\mathcal{P}$, written as:
@@
which remains the top-$k$ neighbors of all locations. The spatial proximity function is usually implemented as Euclidean distance for simplicity.
Temporal Transition Graph. The SDG can reveal the spatial proximity of locations, but it ignores mobility patterns of human trajectories. Consequently, we further propose the TTG to encode mobility patterns based on the partially observed human trajectories. Let $s=[\cdots,\tau_{l},\tau_{l+1},\cdots]$ be a human trajectory and $\tau_{l}=(p_{i},t_{l}),\tau_{l+1}=(p_{j},t_{l+1})$. We record the proximity score between $p_{i}$ and $p_{j}$ as $\epsilon(p_{i},p_{j})=1$ and obtain the summed scores by $w_{ij}=\sum_{(pi,p_{j})\in\mathcal{S}}\epsilon(p_{i},p_{j})$. Then the TTG can be formulated as:
@@
The obtained TTG can provide detailed descriptions of mobility patterns from the temporal transition perspective.
SpatioTemporal Graph. The SDG and TTG describe the relationships between locations from the spatial and temporal perspectives, respectively, ignoring the other aspect to some extent. Therefore, we further characterize the functionality of locations with the STG by their visit distribution over time, which can describe the spatial proximity by spatial cooperation effects and the temporal proximity by the similarity in visit distribution. For instance, cafes often open from morning to evening, while bars keep open till midnight. The visit distribution $F_{p_{i}}$ of a location $p_{i}$ is its normalized visit counts in the observed records $\mathcal{S}_{p_{i}}$ by discretizing the timestamps into $T$ time slots: $F_{p_{i}}(t)=\frac{1}{|\mathcal{S}_{p_{i}}|}\sum_{t_{k}\in\mathcal{S}_{p_{i}}}\mathbbm{1}{[t_{k}=t]}$. The obtained visit distribution $F_{p_{i}}$ can well describe the location $p_{i}$ from both spatial and the temporal aspects. On the one hand, neighbor locations often share similar visit distributions based on geographical effects, which have been widely adopted in Location-based Services. On the other hand, locations far apart can also share similar visit distributions as they provide comparable functionalities like eating or drinking. Thus, the Wasserstein Distance [63, 64] is employed here to quantify the differences among the visit distributions $F_{\mathcal{P}}$ of locations to differentiate their individual functionalities, defined as follows:
@@
where $\pi(x,y)$ is a joint distribution of $F_{p_{i}}$ and $F_{p_{j}}$. The proximity score can be simply defined as $\epsilon(p_{i},p_{j})=1-d(p_{i},p_{j})\in[0,1]$. Let $w_{ij}=\epsilon(p_{i},p_{J})$, the STG can be formulated as:
@@
The proposed STG can capture spatiotemporal dynamics of locations in a distribution way.
Overall, the above multi-channel location graphs are denoted by $\mathcal{G}=\{G_{\mathrm{SDG}},G_{\mathrm{TTG}},G_{\mathrm{STG}}\}$. The connection edges of a location pair $(p_{i},p_{j})\in G$ are naturally weighted with their proximity scores, indicating the spatiotemporal correspondences given by different measurements. However, these scores are not aligned with the human mobility simulation task, which may cause the inconsistent problems between the spatiotemporal graphs and the human mobility simulation task. To bridge the gap between the graphs and the task, we alternatively propose a Vanilla version based on the Weighted version, written as:
@@
The Vanilla graph regards all neighbors as the same, benefiting the model optimization without the need for computed weights of the multi-channel graphs.
@@ 4.1.2 Location Embeddings
Firstly, we implement the features of the location set $\mathcal{P}$ at the input layer by a learnable embedding matrix with end-to-end optimization, following the common training paradigm of human mobility simulation.
Secondly, to adaptively aggregate the useful information in terms of the graph heterogeneity, we adopt the attention layer of graphs [48] to perform multi-channel attention over all kinds of spatiotemporal graphs, namely $\mathcal{G}=\{G_{\mathrm{SDG}},G_{\mathrm{TTG}},G_{\mathrm{STG}}\}$. Given location embeddings $\mathbf{H}^{l-1}_{G}$ from the previous $l-1$-th layer (initialized embeddings at the input layer), our spatiotemporal attention layer computes the node embeddings in a multi-head manner by:
@@
For simplicity, the $l$-th layer location embeddings are summed up over all graphs: $\mathbf{H}^{l}=\sum_{G\in\mathcal{G}}\mathbf{H}^{l}_{G}$. The final embeddings of locations are denoted by $\mathbf{H}_{\mathcal{P}}$.
@@ 4.2 Decision Generator Module
The multi-channel embedding module provides location embeddings via our dedicated spatiotemporal graphs, enhancing the representation quality used in the following decision generator module. The decision generator module autoregressively predicts the next location based on the observed partial mobility trajectory. To cope with the varying durations of locations as shown in Figure 1 (b), we develop the dwell branch besides the common exploration branch in GANs, which also reduces the optimization difficulty to some extent. Since the policy discriminator will not give classification results until the trajectory sequence is finished, we adopt the Monte-Carlo strategy to iteratively predict the next location until the required length.
@@ 4.2.1 Exploration Branch
Suppose that a partial mobility trajectory $s_{1:l}=\{(p_{1},t_{1}),(p_{2},t_{2}),\cdots,(p_{l},t_{l})\}$ is observed, the exploration branch aims to predict the next location $\hat{p}_{l+1}$ that maximizes the decision reward given by the policy discriminator module, where $t_{l+1}$ is assumed uniformly spaced with $t_{l}$ for simplicity. Given a bunch of location embeddings $\mathbf{H}_{\mathcal{P}}$ from the multi-channel embedding module, we firstly fetch the corresponding input embeddings of the trajectory, denoted by $\mathbf{h}_{s}=\{\mathbf{h}_{p_{0}},\mathbf{h}_{p_{1}},\cdots,\mathbf{h}_{p_{l}}\}$. To facilitate the sequence modeling and training efficiency, we adopt the Gated Recurrent Unit [65] as our sequence model, written as:
@@
where the hidden state $z_{l-1}$ is initialized as zero vectors in the first step. The obtained sequence representation $\mathbf{h}_{l}$ is fed into a linear layer to predict the probability of the next locations by:
@@
Then the outputs of the exploration branch $\hat{p}_{l+1}$ together with the outputs of the dwell branch determine the next location iteratively until the sequence is completed.
@@ 4.2.2 Dwell Branch
Motivated by the varying durations of locations in mobility trajectories, we specially design a dwell branch to predict the probability of staying at the previous location, which can inform the model to dwell or explore. Intuitively, the duration of a specific location is determined by the spatiotemporal context and the cumulative duration of a trajectory, where the former can be well captured by the exploration branch and the latter lacks specific modeling. For a highly repetitive trajectory dataset, the optimization goal will be overwhelmed by the repetitive locations, leading to the ignorance of diverse behaviors. Therefore, it is necessary to build the dwell branch to predict whether to stay at the previous location to alleviate the learning difficulties of the exploration branch. Taking the sequence representation $\mathbf{h}_{l}$ for the dwell branch classification is straightforward. However, it ignores the saturation of human mobility behaviors [1] that the durations of locations are usually upper-bounded to some extent. We thus combine the sigmoid function with an exponentially decaying coefficient to reduce the effects of a long-duration location, written as:
@@
Finally, the decision generator module balances the location predictions $\hat{p}_{l+1}$ of the exploration branch and the dwell predictions $\hat{y}_{l+1}^{d}$ by sampling from the output distributions, written as:
@@ 4.3 Policy Discriminator Module
By generating and sampling iteratively in the decision generator module, a set of generated trajectories $\mathcal{T}_{G}$ is fed into the policy discriminator module $D_{\phi}$. Since a powerful discriminator might hinder the optimization of the generator $G_{\theta}$, we build a simple yet efficient discriminator $D_{\phi}$ that contains an embedding matrix of locations and classifies the trajectory sequence as real data or fake data based on a GRU layer, written as:
@@
where $\mathcal{T}_{R}$ and $\mathcal{T}_{G}$ are real and generated trajectories, respectively. The decision generator module $G_{\theta}$ makes stepwise predictions and receives the policy gradients from the discriminator $D_{\phi}$ by unfolding the classification reward recursively following the REINFORCE algorithm [66], written as:
@@ 5 Experiment
TABLE II: Performance of the proposed STAR framework and baselines in terms of JSD for human mobility trajectory simulation. All experimental results are conducted over five trials for a fair comparison. A lower JSD value indicates a better performance. Bold and underline means the best and the second-best results. “*” implies statistical significance for $p<0.05$ under paired t-test.
@@
Markov[10] 0.1253 0.3711 0.0037 0.2686 0.0966 0.0673 0.1373 0.4249 0.0079 0.2627 0.0420 0.0624
IO-HMM[11] 0.4364 0.6170 0.0025 0.0834 0.4053 0.0435 0.3598 0.5514 0.0044 0.1885 0.2325 0.0498
LSTM[67] 0.1240 0.3830 0.0030 0.0862 0.0723 0.0564 0.1249 0.4341 0.0019 0.0426 0.0342 0.0638
DeepMove[12] 0.4292 0.6347 0.0050 0.3113 0.1091 0.0521 0.3518 0.5791 0.0010 0.0480 0.2232 0.0522
GAN[15] 0.4407 0.6272 0.0046 0.0813 0.2420 0.0482 0.3633 0.5615 0.0044 0.1871 0.1679 0.0457
SeqGAN[35] 0.1737 0.4536 0.0028 0.0626 0.0822 0.0563 0.1304 0.4344 0.0011 0.0483 0.0428 0.0571
MoveSim[13] 0.2932 0.5471 0.0021 0.3903 0.1872 0.0574 0.2225 0.4368 0.0054 0.4651 0.1494 0.0510
CGE[14] 0.4098 0.5880 0.0598 0.3189 0.3170 0.0499 0.3913 0.6310 0.0762 0.3984 0.1891 0.0473
VOLUNTEER[68] 0.2498 0.6811 0.6652 0.6931 0.1160 0.0778 0.1804 0.5570 0.6595 0.6931 0.0605 0.0692
STAR (Ours) 0.1134* 0.3615* 0.0012* 0.0597 0.0378* 0.0550 0.1206* 0.4198* 0.0004* 0.0340* 0.0307* 0.0650
@@
Markov[10] 0.0090 0.0589 0.0135 0.1690 0.1231 0.0666 0.0093 0.0389 0.0060 0.1425 0.2517 0.0783
IO-HMM[11] 0.1048 0.0778 0.0096 0.2492 0.2158 0.0724 0.0509 0.0447 0.0041 0.1595 0.3569 0.0577
LSTM[67] 0.0117 0.0194 0.0075 0.1373 0.0358 0.0618 0.0084 0.0111 0.0113 0.1190 0.1186 0.0680
DeepMove[12] 0.1048 0.0854 0.0050 0.1283 0.2552 0.0647 0.0413 0.0338 0.0012 0.0469 0.1410 0.0474
GAN[15] 0.1059 0.0746 0.0096 0.2476 0.2158 0.0714 0.0504 0.0433 0.0043 0.1600 0.3325 0.0577
SeqGAN[35] 0.0222 0.0314 0.0043 0.0615 0.0268 0.0631 0.0145 0.0149 0.0037 0.0924 0.0269 0.0520
MoveSim[13] 0.0828 0.0598 0.0114 0.2739 0.1404 0.0644 0.0222 0.0125 0.0110 0.3473 0.1718 0.0552
CGE[14] 0.0647 0.0981 0.0097 0.2319 0.1968 0.0698 0.0906 0.0494 0.0802 0.4196 0.3286 0.0559
VOLUNTEER[68] 0.0087 0.0186 0.6352 0.6795 0.0843 0.0625 0.0119 0.0112 0.6619 0.6931 0.1137 0.0666
STAR (Ours) 0.0071 0.0165* 0.0022* 0.0516 0.0196* 0.0518* 0.0080 0.0108 0.0009 0.0330* 0.0283 0.0418*